OpenStudy (anonymous):

Simplify 9 a cubed minus 12 a squared plus 6 a all over negative 3 a

OpenStudy (johnweldon1993):

\[\large \frac{9a^3 - 12a^2 + 6a}{-3a}\] Factor out an 'a' from the numerator \[\large \frac{a(9a^2 - 12a + 6)}{-3a}\] Now also notice how every number in the numerator is dividable by 3 (so lets factor our a -3 here) \[\large \frac{-3a(-3a^2 + 4a -2)}{-3a}\] Why did I factor out -3? well \[\large \frac{\cancel{-3a}(-3a^2 + 4a -2)}{\cancel{-3a}}\] So what we're left with is \(\large -3a^2 + 4a - 2\)

OpenStudy (anonymous):

Thanks

OpenStudy (johnweldon1993):

No problem! Did everything make sense? :)

OpenStudy (anonymous):

Yes, i just get confused on theese problems. And never know where to start.